OpenStudy (anonymous):

for each given real number s, find real number t in the interval [0, 2π) so that the point on the unit circle p(t) = p(s) 15π/7 -9π 4π -21π/2 I got the answers but i want to know how to solve these problem, please help :(

OpenStudy (anonymous):

if this is chapter 4 or 5 and I remember right don't you just add the 2pi and continue around the circle?

OpenStudy (anonymous):

If you have a negative point on the unit circle, add 2pi until you get a positive number. Then, for positive values, take out multiples of 2pi For example: \[ \frac{ 15\pi }{ 7} =2\frac{ \pi }{7 } \rightarrow \frac{ \pi }{7}\]